40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15607     Entry time: Fri Oct 2 10:29:49 2020
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: ETMY, BS and ITMX HeNes degrading 

Attachment #1 shows that the ITMX, ETMY and beamsplitter Oplev light levels have decayed significantly from their values when installed. In particular, the ETMY and ITMX sum channels are now only 50% of the values when a new HeNe was installed. ELOG search revealed that ITMY and ETMX HeNes were replaced with newly acquired units in March and September of last year respectively. The ITMX oplev was also replaced in March 2019, but the replacement was a unit that was being used to illuminate our tourist attraction glass fiber at EX.

We should replace these before any vent as they are a useful diagnostic for the DC alignement reference.

Attachment 1: OLsum.png  275 kB  Uploaded Fri Oct 2 11:29:58 2020  | Hide | Hide all
OLsum.png
ELOG V3.1.3-